OpenStudy (anonymous):

solve the equation 3/2x - 5/3x = 2

OpenStudy (lgbasallote):

do you know how to find the LCD/LCM?

OpenStudy (anonymous):

i think so...

OpenStudy (anonymous):

I do actually

OpenStudy (anonymous):

but how is that going to help me @lgbasallote

OpenStudy (lgbasallote):

you'll see.... what is your LCD in this case?

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

or just 6

OpenStudy (lgbasallote):

great! so your LCD is 6x... the fractions would then become \(\large \frac{3(3)}{6x} - \frac{5(2)}{6x} = 2\) agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

now combine your fractions..

OpenStudy (anonymous):

so it would be 9/6x - 10/ 6x =2

OpenStudy (anonymous):

then -1/6x?

OpenStudy (lgbasallote):

right! -1/6x = 2 now cross multiply

OpenStudy (anonymous):

12x/-1?

OpenStudy (lgbasallote):

no no... -1 = 12x

OpenStudy (lgbasallote):

from that you remove the coefficient of x

OpenStudy (anonymous):

so does x = -1/12

OpenStudy (lgbasallote):

\(\Huge \checkmark\)

OpenStudy (anonymous):

thanks for the help

OpenStudy (lgbasallote):

<tips imaginary hat>